# Robust affine point matching via quadratic assignment on Grassmannians

### v4

@@ Abstract
Robust Affine matching with Grassmannians (RAG) is a new algorithm to perform affine registration of point clouds. The algorithm is based on minimizing the Frobenius distance between two elements of the Grassmannian. For this purpose, an indefinite relaxation of the Quadratic Assignment Problem (QAP) is used, and several approaches to affine feature matching are studied and compared. Experiments demonstrate that RAG is more robust to noise and point discrepancy than previous methods.
Index Terms:
@@
I Introduction
@@
II Algorithm RAG
@@
Let $\mathrm{Gr}(k,n)$ be the Grassmannian of $k$–dimensional planes in $\mathbb{R}^{n}$. We can view the point cloud $X$ as a map $X:\mathbb{R}^{n}\rightarrow\mathbb{R}^{d}$ defined by $X(v)=X\,v$, for $v\in\mathbb{R}^{n}$. Then $\ker X$ is an element of $\mathrm{G}=\mathrm{Gr}(n-d,n)\cong\mathrm{Gr}(d,n)$. Moreover, for any $L\in\mathrm{GL}(d)$ we have $\ker LX=\ker X$. Thus, $X$ and all of its non–degenerate images define the same element of $\mathrm{G}$.
Let $P_{X}$ be the orthogonal projection of $\mathbb{R}^{n}$ onto $\ker X^{\perp}$. The orthogonal projection $P_{X}$, which uniquely defines $\ker X$, will be used as the representation of an element of $\mathrm{G}$. Note that $P_{X}$ can be computed from the (reduced) singular value decomposition (SVD) of $X$,
@@
Thus, we can formulate our problem as
@@
Relaxing the feasible region allows one to employ the tools of continuous optimization to search for local optima. As this is not a convex problem different initializations can give different results.
@@
where $C>0$ is some large constant. The requirement of $w(S)$ is that $w(S)\approx 1$ for permutations $S$ that are Hamming–close to $S_{*}$ and $w(S)\approx 0$ for all others.
@@
Another approach would be to take such an $S_{*}$ among $S_{1},\ldots,S_{N}$ that provides the best match between $P_{X}$ and $P_{Y}$, that is $S_{*}=\textrm{argmin}_{S_{i}}\|P_{Y}-S_{i}^{T}P_{X}S_{i}\|_{F}$. The results of [8] suggest that this indefinite relaxation (rQAP) of QAP proposed in [10] finds the optimal solution, $S_{*}$, with high probability.
We compare the best match and weighted sum approaches in section  II-D below.
@@
Algorithm RAG
@@
for $i\leftarrow 1$ to $N$ do $S_{i}\leftarrow\max_{S\in\mathrm{Sym}(n)}\mathrm{tr}(P_{Y}\,S^{T}\,P_{X}\,S)$triangleright rQAP
@@
II-A Correctness
Let $S$ be the feature matching matrix produced by RAG. Let $\mathbb{P}(S=S_{*})=1-\alpha$ for some constant $10\leq\alpha<$ 1, and $\mathbb{P}(S\neq S_{*})=\alpha$. In the setting of best matching for random Bernoulli graphs $\alpha=0$ holds [10].
@@ II-A Correctness
Thus, once $\varepsilon<\frac{1-\alpha}{\alpha}$ we have that $\pi(\mathbb{E}\,S_{w})=S_{*}$. However, we need to guarantee that the same equality holds once we replace $\mathbb{E}\,S_{w}$ by the sample average $\overline{S}_{w}$ for a large enough sample $\{S_{i}\}^{N}_{i=1}$ of permutation matrices produced by RAG.
@@ II-C Point discrepancy
Then we replace the projection matrix $P_{X}\in\mathbb{R}^{m\times m}$ with $P_{X}\leftarrow P_{X}\oplus\mathbf{0}_{n-m}$ and apply RAG to the new pair $P_{X},P_{Y}\in\mathbb{R}^{n\times n}$. Once an output $S_{0}$ is produced, we pick only the first $m$ columns to produce a matching between the points of $X$ and $Y$. Indeed, matching any of the extra $n-m$ rows/columns of $P_{Y}$ to a zero row/column in $P_{X}$ will be heavily penalized, and thus rQAP will target the closest possible match for the first $m$ rows/columns of $P_{X}$ and some $m$ rows/columns of $P_{Y}$. Based on our experiments, this approach works well with missing/added points in $X$ and $Y$, even for relatively large discrepancies.
II-D Finding feature matching
@@
As evidenced in Figure 17 nether approach shows acceptable accuracy for $N\leq 2^{8}$. For $N\geq 2^{9}$ the best match approach shows equally good or better accuracy while the noise is relatively small ($0.150.0\sigma<$ 0.15), while the weighted sum approach appears to be more robust for noise of larger magnitudes ($0.250.15<\sigma<$ 0.25).
III Experiments
@@
III-A Random point clouds
In our numerical experiments, RAG was applied to pairs of synthetic point clouds with the following specifications: one specimen point cloud $X\subset\mathbb{R}^{3\times m}$ and another point cloud $Y\subset\mathbb{R}^{3\times n}$ such that
@@ III-A Random point clouds
III-B Caerbannog point clouds
@@
For each value of the noise $\sigma$ and point discrepancy $\lambda$, a batch of $10$ tests was run and the mean values of $\delta_{L}$, $\delta_{Y}$, and $\delta_{X}$ were computed. Every run of RAG used $N=2^{10}$ calls to rFAQ.
The Python code to perform the tests and the respective output stored is accessible on GitHub [14]. The code requires SageMath [15] and runs in the Jupiter environment. The output is saved in csv format and can be accessed without the need to run the code first.
@@
    $d_{\sigma}=\|Y^{\prime}-Y\|_{2}/\|Y\|_{2}$ (the relative noise as compared to the ground truth image of $X$ under $L$); and
@@
    $d_{\lambda}=\|X^{\prime}\setminus X\|_{2}/\|X\|_{2}$ (the relative point discrepancy as compared to the initial specimen image).
@@
In Figures 4, 4 and 4, we show the specimen cloud $X$ (yellow), the preimage $[L^{-1}\,Y\,S^{T}]_{X}$ (red), and also the point discrepancy consisting of points in $X^{\prime}$ that do not belong to $X$ (blue). For all three point clouds we have $\sigma=0.05$, $\lambda=0.90$, $\mathrm{cond}\,L=3$. The number of trials in each case is $N=2^{10}$.
@@
III-C Partial scans of the same object
@@
In Figure 14 (center) the ground truth of matching the damaged and restored sculptures is shown. In Figure 1516, we show the matching of the broken sculpture (from which $X$ is sampled) to the restored one with a linear transform $L$ applied (from which $Y$ is sampled). In order not to produce a confusing image of the sculpture distorted by $L$, we instead produce the image of the restored sculpture under $L^{-1}_{0}L$ (where $L$ is the original linear transform, and $L_{0}$ is recovered by running RAG) followed by a translation. The translation adjusts the barycenters of the two meshes.
As compared to rigid registration, affine registration cannot recover the initial map so well in the presence of occlusion and noise: the affine group has more degrees of freedom, and thus it will necessarily register some bias produced by the noise and occlusion points. Note that in Figure 15 with more sample points and a smaller condition number of the random linear map, we have a better match than in Figure 16 where the number of sample points is reduced and the condition number of the random linear map is larger.
The code and mesh data used are available on GitHub [14].
IV Comparison to other algorithms
In this section, we compare RAG to another algorithm, GrassGraf [7], that appears to be the state–of–the–art algorithm to recover affine correspondences.
@@ IV Comparison to other algorithms
The main finding is that GrassGraf loses precision very quickly if $Y$ is affected even by relatively small noise. The test point clouds that were used are the unoccluded Caerbannog clouds [13]. We measured the values $\delta_{L}$ and $\delta_{Y}$, just as in the case of RAG. These values are already enough to conclude that GrassGraph does not appear to recover $L$ robustly.
In Figures 2323, we show the graphical depiction of our test results (as in the case of RAG) for GrassGraph with $3$ Laplacian eigenvectors used. The analogous dataset with test results for GrassGraph with $10$ Laplacian eigenvectors used is available on GitHub [14]. We do not reproduce it here as it appears to show only a marginal improvement.
Also, some images of the test clouds are available on GitHub [14]: we do not reproduce them here, just note that the “Teapot” cloud shows slightly better robustness to noise and point discrepancy while, say, the “Cow” cloud already demonstrates a great difference between the recovered $L_{0}$ and the initial $L$ even for relatively small noise and point discrepancy.
V Conclusions
@@
This paper provides a robust and efficient algorithm to find correspondences between point clouds up to an affine transformation and gives bounds on the robustness of the algorithm in the presence of noise and point discrepancies.
@@
VI Acknowledgements
The authors are grateful to the Department of Computer Science, University of Neuchâtel, for the opportunity to use their computational cluster Phi–2 for the numerical experiments. This work was supported by the Swiss National Science Foundation [grant no. PP00P2–202667] and by an Israel Science Foundation (ISF) grant.

### v5

@@ Abstract
Robust Affine Matching with Grassmannians (RoAM) is a new algorithm to perform affine registration of point clouds. The algorithm is based on minimizing the Frobenius distance between two elements of the Grassmannian. For this purpose, an indefinite relaxation of the Quadratic Assignment Problem (QAP) is used, and several approaches to affine feature matching are studied and compared. Experiments demonstrate that RoAM is more robust to noise and point discrepancy than previous methods.
keywords:
@@
1 Introduction
Figure 2: “Teapot” cloud: specimen $X$ (yellow) and preimage $L^{-1}_{0}\,Y$ (red). Points of $X$ having no correspondence in $Y$ are marked (blue).
Figure 3: “Bunny” cloud: specimen $X$ (yellow) and preimage $L^{-1}_{0}\,Y$ (red). Points of $X$ having no correspondence in $Y$ are marked (blue).
Figure 4: “Cow” cloud: specimen $X$ (yellow) and preimage $L^{-1}_{0}\,Y$ (red). Points of $X$ having no correspondence in $Y$ are marked (blue).
@@ 1 Introduction
2 Algorithm RoAM
The algorithm is based on two main ideas, first affine invariant representations, Grassmanians, are computed for the point clouds and then a relaxation of a hard computational problem (QAP) is used to find the matching between the representative Grassmanians.
@@ 2 Algorithm RoAM
Let $\mathrm{Gr}(k,n)$ be the Grassmannian of $k$–dimensional planes in $\mathbb{R}^{n}$. We can view the point cloud $X$ as a map $X:\mathbb{R}^{n}\rightarrow\mathbb{R}^{d}$ defined by $X(v)=X\,v$, for $v\in\mathbb{R}^{n}$. Then $\ker X$ is an element of $\mathrm{G}=\mathrm{Gr}(n-d,n)\cong\mathrm{Gr}(d,n)$. Moreover, for any $L\in\mathrm{GL}(d)$ we have $\ker LX=\ker X$. Thus, $X$ and all of its non–degenerate images define the same element of $\mathrm{G}$. Let $P_{X}$ be the orthogonal projection of $\mathbb{R}^{n}$ onto $\ker X^{\perp}$. The orthogonal projection $P_{X}$, which uniquely defines $\ker X$, will be used to represent an element of $\mathrm{G}$. Note that $P_{X}$ can be computed from the (reduced) singular value decomposition (SVD) of $X$,
@@
Thus, we can reformulate the initial problem as
@@
Relaxing the feasible region allows one to employ the tools of continuous optimization to search for local optima, as this is not a convex problem different initializations can give different results.
@@
The penultimate equality holds because $P^{2}=P$ for any projection operator $P$.
@@
where $C>0$ is some large constant. The requirement is that $w(S)\approx 1$ for permutations $S$ that are Hamming–close to $S_{*}$ and $w(S)\approx 0$ for all others.
@@
Another approach would be to take such an $S_{*}$ among $S_{1},\ldots,S_{N}$ that provides the best match between $P_{X}$ and $P_{Y}$, that is $S_{*}=\textrm{argmin}_{S_{i}}\|P_{Y}-S_{i}^{T}P_{X}S_{i}\|_{F}$. The results of [8] suggest that this indefinite relaxation of QAP proposed in [10] finds the optimal solution, $S_{*}$, with high probability. We shall refer to the algorithm proposed in [8] as rQAP (relaxed Quadratic Assignment Problem).
We compare the best match and weighted sum approaches in section  2.4 below.
@@
Algorithm 1 RoAM: Robust Affine Matching with Grassmannians
@@
for $i\leftarrow 1$ to $N$ do $S_{i}\leftarrow\max_{S\in\mathcal{D}}\mathrm{tr}(P_{Y}\,S^{T}\,P_{X}\,S)$triangleright rQAP
@@
2.1 Correctness
Let $S$ be the feature matching matrix produced by RoAM. Let $\mathbb{P}(S=S_{*})=1-\alpha$ for some constant $10\leq\alpha<$ 1, and $\mathbb{P}(S\neq S_{*})=\alpha$. In the setting of best matching for random Bernoulli graphs $\alpha=0$ holds [10].
@@ 2.1 Correctness
Thus, once $\varepsilon<\frac{1-\alpha}{\alpha}$ we have that $\pi(\mathbb{E}\,S_{w})=S_{*}$. However, we need to guarantee that the same equality holds once we replace $\mathbb{E}\,S_{w}$ by the sample average $\overline{S}_{w}$ for a large enough sample $\{S_{i}\}^{N}_{i=1}$ of permutation matrices produced by RoAM.
@@ 2.3 Point discrepancy
Then we replace the projection matrix $P_{X}\in\mathbb{R}^{m\times m}$ with $P_{X}\leftarrow P_{X}\oplus\mathbf{0}_{n-m}$ and apply RoAM to the new pair $P_{X},P_{Y}\in\mathbb{R}^{n\times n}$. Once an output $S_{0}$ is produced, we pick only the first $m$ columns to produce a matching between the points of $X$ and $Y$. Indeed, matching any of the extra $n-m$ rows/columns of $P_{Y}$ to a zero row/column in $P_{X}$ will be heavily penalized, and thus rQAP will target the closest possible match for the first $m$ rows/columns of $P_{X}$ and some $m$ rows/columns of $P_{Y}$. Based on our experiments, this approach works well with missing/added points in $X$ and $Y$, even for relatively large discrepancies.
2.4 Finding feature matching
@@
As evidenced in Figure 19 neither approach shows acceptable accuracy for $N\leq 2^{8}$. For $N\geq 2^{9}$ the best match approach shows equally good or better accuracy while the noise is relatively small ($0.150.0<\sigma<$ 0.15), while the weighted sum approach appears to be more robust for noise of larger magnitudes ($0.250.15<\sigma<$ 0.25).
3 Experiments
@@
3.1 Random point clouds
In our numerical experiments, RoAM was applied to pairs of synthetic point clouds with the following specifications: one specimen point cloud $X\subset\mathbb{R}^{3\times m}$ and another point cloud $Y\subset\mathbb{R}^{3\times n}$ such that
  • 1.
@@ 3.1 Random point clouds
3.2 Caerbannog point clouds
@@
For each value of the noise $\sigma$ and point discrepancy $\lambda$, a batch of $10$ tests was run and the mean values of $\delta_{L}$, $\delta_{Y}$, and $\delta_{X}$ were computed. Every run of RoAM used $N=2^{10}$ calls to rQAP.
The Python code to perform the tests and the respective output stored is accessible on GitHub [14]. The code requires SageMath [15] and runs in the Jupiter environment. The output is saved in csv format and can be accessed without running the code.
For different levels of noise and occlusion, we got $69.43$ seconds on average per batch of $100$ trials for the “choose best” approach and $57.07$ seconds on average per batch of $100$ trials for the “weighted sum” approach (see [14]). Thus, a single run of RoAM took approximately between $0.57$ to $0.69$ seconds on a MacBook Pro (M1, 2020).
@@
    $d_{\sigma}=\|Y^{\prime}-Y\|_{2}/\|Y\|_{2}$ (the relative noise compared to the ground truth image of $X$ under $L$); and
@@
    $d_{\lambda}=\|X^{\prime}\setminus X\|_{2}/\|X\|_{2}$ (the relative point discrepancy compared to the initial specimen image).
@@
In Figures 2, 3 and 4, we show the specimen cloud $X$ (yellow), the preimage $[L^{-1}\,Y\,S^{T}]_{X}$ (red), and also the point discrepancy consisting of points in $X^{\prime}$ that do not belong to $X$ (blue). For all three point clouds we have $\sigma=0.05$, $\lambda=0.90$, $\mathrm{cond}\,L=3$. The number of trials in each case is $N=2^{10}$.
@@
3.3 Partial scans of the same object
@@
In Figure 14 (center) the ground truth of matching the damaged and restored sculptures is shown. In Figure 1516, we show the matching of the broken sculpture (from which $X$ is sampled) to the restored one with a linear transform $L$ applied (from which $Y$ is sampled). In order not to produce a confusing image of the sculpture distorted by $L$, we instead show the image of the restored sculpture under $L^{-1}_{0}L$ (where $L$ is the original linear transform, and $L_{0}$ is recovered by running RoAM) followed by a translation. The translation adjusts the barycenters of the two meshes.
Affine registration cannot recover the initial map as well as rigid registration in the presence of occlusion and noise: the affine group has more degrees of freedom, and thus it will necessarily register some bias produced by the noise and occlusion points. Note that in Figure 15 with more sample points and a smaller condition number of the random linear map, we have a better match than in Figure 16 where the number of sample points is reduced and the condition number of the random linear map is larger.
The code and mesh data are available on GitHub [14].
4 Comparison to other algorithms
In this section, we compare RoAM to another algorithm, GrassGraf [7], which appears to be the state–of–the–art algorithm to recover affine correspondences.
@@ 4 Comparison to other algorithms
The main finding is that GrassGraf loses precision very quickly if $Y$ is affected even by relatively small noise. The test point clouds used are the unoccluded Caerbannog clouds [13]. We measured the values $\delta_{L}$ and $\delta_{Y}$, just as in the case of RoAM. These values are enough to conclude that GrassGraph does not appear to recover $L$ robustly.
In Figures 2118, we show the graphical depiction of our test results (as in the case of RoAM) for GrassGraph with $3$ Laplacian eigenvectors used. The analogous dataset with test results for GrassGraph with $10$ Laplacian eigenvectors used is available on GitHub [14]. We do not reproduce it here as it shows only a marginal improvement.
Also, some images of the test clouds are available on GitHub [14]: we do not reproduce them here, note that the “Teapot” cloud shows slightly better robustness to noise and point discrepancy while, say, the “Cow” cloud already demonstrates a great difference between the recovered $L_{0}$ and the initial $L$ even for relatively small noise and point discrepancy.
5 Conclusions
@@
This paper provides a robust and efficient algorithm to find correspondences between point clouds up to an affine transformation and gives bounds on the algorithm’s robustness in the presence of noise and point discrepancies.
@@
6 Acknowledgements
This research is supported by the University of Austin (UATX) and by the Israel Science Foundation (ISF).